Two hot dogs and one hamburger cost $3.39. Three hot dogs and three hamburgers cost $6.57. What is the cost of one hot dog?
Answer by LilSkittleMd(119)

Let d=hotdogs; h=hamburger
The two equations are
2d+h=3.39
3d+3h=6.57
You can use the substitution method to solve this. Take the first equation and solve for h. The equation for that is:
h=-2d+3.39
Susbstitute that into the second equation
3d+3(-2d+3.39)=6.57
3d-6d+10.17=6.57
-3d=-3.6
d=1.2
The cost of one hotdog is $1.20